# The Complexity of Sequential Prediction in Dynamical Systems

### v1

@@ 2.4 Combinatorial dimensions
In addition to complexity measures, combinatorial dimensions play an important role in providing crisp quantitative characterizations of learnability. For example, the Daniely Shalev-Shwartz dimension (DSdim), originally proposed by Daniely and Shalev-Shwartz (2014) and formally defined in Appendix A , was recently shown by Brukhim et al. (2022) to provide a tight quantitative characterization of multiclass PAC learnability. In Section 3.3, we use the DSdim to relate the realizable learnability of dynamical systems to multiclass PAC learnability of $\mathcal{F}$.
Analogously, for online multiclass classification, the Littlestone dimension (Ldim), originally proposed by Littlestone (1987) for binary classification and later extended to multiclass classification by Daniely et al. (2011), provides a tight quantitative characterization of learnability (Hanneke et al., 2023). We also define the Littlestone dimension in Appendix A.
In this tradition, we are interested in identifying combinatorial dimensions that provide both qualitative and quantitative characterizations of the learnability of dynamical systems. In Section 4, we show that the Littlestone dimension of $\mathcal{F}$ characterizes agnostic learnability of dynamical systems. On the other hand, in Section 3, we show that a new scale-sensitive dimension, termed the Evolution dimension, characterizes the realizable learnability of dynamical systems.
Definition 6 (Evolution dimension).
The evolution dimension of $\mathcal{F}$ at scale $\gamma>0$, denoted $\emph{{E}}_{\gamma}(\mathcal{F})$, is the largest $d\in\mathbb{N}$ such that there exists a trajectory tree $\mathcal{T}$ of depth $d$ shattered by $\mathcal{F}$ with $\emph{{B}}(\mathcal{T})\geq\gamma d$. If this is true arbitrarily large depths $d\in\mathbb{N}$, we say $\emph{{E}}_{\gamma}(\mathcal{F})=\infty$.
3 Realizable Learnability
3.1 Characterizations of Realizable Learnability
In this section, we provide qualitative and quantitative characterizations of realizable learnability in terms of the Evolution complexity and dimension. Our main result in this section is Theorem 3.1, which provides bounds on the minimax expected number of mistakes.
@@ 3.1 Characterizations of Realizable Learnability
  • (i)
    $\frac{1}{2}\,\emph{{C}}_{T}(\mathcal{F})\leq\inf_{\mathcal{A}}\,\emph{{M}}_{\mathcal{A}}(T,\mathcal{F})\leq\emph{{C}}_{T}(\mathcal{F}).$
  • (ii)
    $\sup_{\gamma>0}\Bigl{\{}\gamma\emph{{E}}_{\gamma}(\mathcal{F})-\max\{\emph{{E}}_{\gamma}(\mathcal{F})-T,0\}\Bigl{\}}\leq\emph{{C}}_{T}(\mathcal{F})\leq\inf_{\gamma>0}\{\gamma\max\{\emph{{E}}_{\gamma}(\mathcal{F}),T\}+\gamma\}.$
Moreover, the upperbound in (i) is achieved constructively by a deterministic learner.
The factor of $\frac{1}{2}$ in the lowerbound of part (i) is due to randomized learners, and one can get $\inf_{\text{Deterministic }\mathcal{A}}\texttt{M}_{\mathcal{A}}(T,\mathcal{F})=\texttt{C}_{T}(\mathcal{F})$. In (ii), for all $\gamma>0$ such that $\texttt{E}_{\gamma}(\mathcal{F})\leq T$, we obtain the bound $\gamma\texttt{E}_{\gamma}(\mathcal{F})\leq\texttt{C}_{T}(\mathcal{F})$. We prove parts (i) of Theorem 3.1 using constructive arguments below. Part (ii) is proven via non-constructive arguments in Appendix B.
Proof.
@@
(of lowerbound in (i) of Theorem 3.1). Let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ be any evolution class and $T\in\mathbb{N}$ be the time horizon. Let $\mathcal{A}$ be any randomized learner. Our goal will be to construct a hard realizable trajectory $\{x_{t}\}_{t=0}^{T}$ such that $\mathcal{A}$’s expected number of mistakes is at least $\frac{\texttt{C}_{T}(\mathcal{F})}{2}$. Without loss of generality, suppose $d:=\texttt{C}_{T}(\mathcal{F})>0$ as otherwise the lowerbound holds trivially. Then, by definition of the Evolution complexity, there exists a trajectory tree $\mathcal{T}$ of depth $T$ shattered by $\mathcal{F}$ with branching factor at least $d$. This means that for every path $\sigma\in\{-1,1\}^{T}$ down $\mathcal{T}$, we have $\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{T}_{t}((\sigma_{<t},-1))\neq\mathcal{T}_{t}((\sigma_{<t},+1))\}\geq d$.
@@
where the first inequality follows from the fact that $\sigma_{t}\sim\{-1,1\}$. ∎
Proof.
(of upperbound in (i) of Theorem 3.1). Let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ be any evolution function class and let $T\in\mathbb{N}$ be the time horizon. Our goal will be to construct a deterministic learner $\mathcal{A}$ such that for any realizable trajectory $\{x_{t}\}_{t=0}^{T}$, $\mathcal{A}$ makes at most $\texttt{C}_{T}(\mathcal{F})$ mistakes. To that end, it will be useful to define an instance-dependent version of the evolution complexity.
@@
$\texttt{C}_{T}(V,x_{0}):=\sup\{\texttt{B}(\mathcal{T})\,\mid\,\mathcal{T}\in\mathcal{S}(V,T)\text{ and }\mathcal{T}\text{ rooted at }x_{0}\}$
denote an instance-dependent Evolution complexity of $V$, where $\mathcal{S}(V,T)$ is set of all trajectory trees of depth $T$ shattered by $V$. Note that $\texttt{C}_{T}(V)=\sup_{x_{0}\in\mathcal{X}}\texttt{C}_{T}(V,x_{0})$. $\mathcal{A}$ is a version-space algorithm that uses the instance-dependent Evolution complexity to make its prediction such that whenever $\mathcal{A}$ errs, the instance-dependent Evolution complexity decreases. In this way, $\texttt{C}_{T}(V,x_{0})$ acts as a potential function. Algorithm 1 formalizes this idea.
@@
Let $V_{1}=\mathcal{F}$ for $t=1,...,T$ do
       For $x\in\mathcal{X}$, define $V_{t}^{x}=\{f\in V_{t}:f(x_{t-1})=x\}.$ if $\{f(x_{t-1}):f\in V_{t}\}=\{x\}$ then
             Predict $\hat{x}_{t}=x$.
       else
             Predict $\hat{x}_{t}\in\operatorname*{arg\,max}_{x\in\mathcal{X}}\texttt{C}_{T-t}(V_{t}^{x},x).$
       Receive $x_{t}$ and update $V_{t+1}\leftarrow V_{t}^{x_{t}}.$
end for
@@
We now show that Algorithm 1 makes at most $\texttt{C}_{T}(\mathcal{F})$ mistakes on any realizable trajectory. Let $\{x_{t}\}_{t=0}^{T}$ denote the realizable trajectory to be observed. It suffices to show that
$\texttt{C}_{T-t}(V_{t+1},x_{t})\leq\texttt{C}_{T-t+1}(V_{t},x_{t-1})-\mathbbm{1}\{x_{t}\neq\hat{x}_{t}\}$ (2)
@@
$\sum_{t=1}^{T}\texttt{C}_{T-t}(V_{t+1},x_{t})\leq\sum_{t=1}^{T}\texttt{C}_{T-t+1}(V_{t},x_{t-1})-\sum_{t=1}^{T}\mathbbm{1}\{x_{t}\neq\hat{x}_{t}\}.$
@@
$\sum_{t=1}^{T}\mathbbm{1}\{x_{t}\neq\hat{x}_{t}\}$ $\leq\sum_{t=1}^{T}\texttt{C}_{T-t+1}(V_{t},x_{t-1})-\sum_{t=1}^{T}\texttt{C}_{T-t}(V_{t+1},x_{t})$
$=\sum_{t=1}^{T}\Bigl{(}\texttt{C}_{T-t+1}(V_{t},x_{t-1})-\texttt{C}_{T-t}(V_{t+1},x_{t})\Bigl{)}$
$=\texttt{C}_{T}(V_{1},x_{0})-\texttt{C}_{0}(V_{T+1},x_{T})\leq\texttt{C}_{T}(\mathcal{F}).$
as needed. To prove (2), we need to consider three cases: (a) $\texttt{C}_{T-t+1}(V_{t},x_{t-1})>0$ and $\hat{x}_{t}\neq x_{t}$, (b) $\texttt{C}_{T-t+1}(V_{t},x_{t-1})>0$ and $\hat{x}_{t}=x_{t}$, and (c) $\texttt{C}_{T-t+1}(V_{t},x_{t-1})=0$.

### v2

@@ 2.4 Combinatorial dimensions
In addition to complexity measures, combinatorial dimensions play an important role in providing crisp quantitative characterizations of learnability. For example, the Daniely Shalev-Shwartz dimension (DSdim), originally proposed by Daniely and Shalev-Shwartz (2014) and formally defined below, was recently shown by Brukhim et al. (2022) to provide a tight quantitative characterization of multiclass PAC learnability. In Section 3.2, we use the DSdim to relate the realizable learnability of dynamical systems to multiclass PAC learnability of $\mathcal{F}$.
Definition 2.7 (DS dimension (Daniely and Shalev-Shwartz, 2014)).
We say that $A\subseteq\mathcal{X}$ is DS-shattered by $\mathcal{F}$ if there exists a finite $\mathcal{H}\subset\mathcal{F}$ such that for every $x\in A$ and $h\in\mathcal{H}$, there exists a $g\in\mathcal{H}$ such that $g(x)\neq h(x)$ and $g(z)=h(z)$ for all $z\in A\setminus\{x\}$. The DS dimension of $\mathcal{F}$, denoted $\operatorname{DS}(\mathcal{F})$, is the largest $d\in\mathbb{N}$ such that there exists a shattered set $A\subset\mathcal{X}$ with cardinality $d$. If there are arbitrarily large sets $A\subseteq\mathcal{X}$ that are shattered by $\mathcal{F}$, then we say that $\operatorname{DS}(\mathcal{F})=\infty.$
Analogously, for online multiclass classification, the Littlestone dimension (Ldim), originally proposed by Littlestone (1987) for binary classification and later extended to multiclass classification by Daniely et al. (2011), provides a tight quantitative characterization of learnability (Hanneke et al., 2023).
Definition 2.8 (Littlestone dimension (Littlestone, 1987; Daniely et al., 2011)).
Let $\mathcal{T}$ be a complete binary tree of depth $d$ whose internal nodes are labeled by a sequence $(\mathcal{T}_{0},\ldots,\mathcal{T}_{d-1})$ of node-labeling functions $\mathcal{T}_{t-1}:\{-1,1\}^{t-1}\to\mathcal{X}$. The tree $\mathcal{T}$ is shattered by $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ if there exists a sequence $(Y_{1},...,Y_{d})$ of edge-labeling functions $Y_{t}:\{-1,1\}^{t}\rightarrow\mathcal{X}$ such that for every path $\sigma=(\sigma_{1},...,\sigma_{d})\in\{-1,1\}^{d}$, there exists a function $f_{\sigma}\in\mathcal{F}$ such that for all $t\in[d]$, $f_{\sigma}(\mathcal{T}_{t-1}(\sigma_{<t}))=Y_{t}(\sigma_{\leq t})$ and $Y_{t}((\sigma_{<t},-1))\neq Y_{t}((\sigma_{<t},+1))$. The Littlestone dimension of $\mathcal{F}$, denoted $\operatorname{L}(\mathcal{F})$, is the maximal depth of a tree $\mathcal{T}$ that is shattered by $\mathcal{F}$. If there exists shattered trees of arbitrarily large depth, we say $\operatorname{L}(\mathcal{F})=\infty$.
3 Warmup: Realizable Learnability
In this section, we provide qualitative and quantitative characterizations of realizable learnability in terms of the Evolution complexity. Our main result in this section is Theorem 3.1, which provides bounds on the minimax expected number of mistakes.
@@ 3 Warmup: Realizable Learnability
For any $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}},$ we have $\frac{1}{2}\,\operatorname{C}_{T}(\mathcal{F})\leq\inf_{\mathcal{A}}\,\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\leq\operatorname{C}_{T}(\mathcal{F}).$ Moreover, the upper bound is achieved constructively by a deterministic learner.
The factor of $\frac{1}{2}$ in the lower bound is due to randomized learners, and the lower bound of $\operatorname{C}_{T}(\mathcal{F})$ can be obtained if the learner is restricted only to deterministic learning rules.
We now describe our high-level proof strategy and defer the full proof of Theorem 3.1 to Appendix A. Our lower bound involves picking the worst-case shattered tree with the largest branching factor and traversing down this tree uniformly at random to generate the sequence of states. For such a sequence, the learner can do no better than random guessing at nodes where the branching occurred, yielding the lower bound $\operatorname{C}_{T}(\mathcal{F})/2$. Next, for our upper bound, we first define a localized complexity measure $\operatorname{C}_{T}(\mathcal{F},x_{0})$, where we only consider shattered trees rooted at the revealed initial state $x_{0}$. Our minimax learner is then a version space algorithm that predicts the state that will result in the largest reduction in the complexity measure if a mistake occurs. This learner is a generalization of the celebrated Standard Optimal Algorithm due to Littlestone (1987).
3.1 Minimax Rates in the Realizable Setting
While Theorem 3.1 provides a quantitative and qualitative characterization of realizable learnability, it does not shed light on how $\inf_{\mathcal{A}}\operatorname{M}_{\mathcal{A}}(\mathcal{F},T)$ may depend on the time horizon $T$. In online classification with the 0-1 loss, the seminal work by Littlestone (1987) and Daniely et al. (2011) show that only two rates are possible: $\Theta(T)$ and $\Theta(1)$. That is, if a hypothesis class is online learnable in the realizable setting, then it is learnable with a constant mistake bound (i.e. the Littlestone dimension). Perhaps surprisingly, this is not the case for learning dynamical systems in a strong sense: every rate is possible.
Theorem 3.2.
For every $S\subset\mathbb{N}\cup\{0\}$, there exists $\mathcal{F}_{S}\subseteq\mathcal{\mathbbm{Z}}^{\mathbbm{Z}}$ such that $\operatorname{C}_{T}(\mathcal{F}_{S})=\sup_{n\in\mathbb{N}\cup\{0\}}|S\cap\{n,n+1,...,n+T-1\}|.$
Theorem 3.2, proved in Appendix B, along with Theorem 3.1 implies that any minimax rate in the realizable setting is possible. As an example, suppose we would like to achieve the rate $\Theta(T^{\alpha})$ for some $\alpha<1$. Then, picking $S=\{\lfloor t^{\frac{1}{\alpha}}\rfloor:t\in\mathbb{N}\cup\{0\}\}$ suffices since $\sup_{x\in\mathbb{N}\cup\{0\}}|S\cap\{n,n+1,...,n+T-1\}|=|\{\lfloor t^{\frac{1}{\alpha}}\rfloor:t\in\mathbb{N}\cup\{0\}\}\cap\{0,1,...,T-1\}|=\Theta(T^{\alpha})$. Likewise, one can get logarithmic rates by picking $S=\{2^{t}:t\in\mathbb{N}\cup\{0\}\}$ and constant rates by picking $S\subset\mathbb{N}\cup\{0\}$ such that $|S|<\infty$. In light of Theorem 3.2 and the fact that only constant mistake bounds are possible for online multiclass classification, it is natural to ask when one can achieve constant mistake bounds for learning dynamical systems. To answer this question, we introduce a new combinatorial dimension termed the Branching dimension.
Definition 3.3 (Branching dimension).
The Branching dimension, denoted $\operatorname{Bd}(\mathcal{F})$, is the smallest natural number $d\in\mathbb{N}$ such that for every shattered trajectory tree $\mathcal{T}$, we have $\operatorname{B}(\mathcal{T})\leq d$. If for every $d\in\mathbb{N}$, there exists a shattered trajectory tree $\mathcal{T}$ with $\operatorname{B}(\mathcal{T})>d$, we say $\operatorname{Bd}(\mathcal{F})=\infty$.
Theorem 3.4, proved via non-constructive arguments in Appendix C, shows that $\inf_{\mathcal{A}}\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})=\Theta(1)$ if and and only if $\operatorname{Bd}(\mathcal{F})<\infty$.
Theorem 3.4 (Constant Minimax Expected Mistakes).
For any $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$, we have (i) $\inf_{\mathcal{A}}\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\leq\operatorname{Bd}(\mathcal{F})$ and (ii) if $\operatorname{Bd}(\mathcal{F})=\infty$, then $\inf_{\mathcal{A}}\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})=\omega(1).$ [^2] Recall that $f(n)=\omega(1)$ if for all $k>0$, there exists $n_{k}$, such that for all $n>n_{k}$ we have $f(n)>k$.
3.2 Relations to PAC and Online Multiclass Classification
By studying abstract state spaces $\mathcal{X}$ and evolutions function classes $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$, we can also compare realizable learnability of dynamical systems to existing notions of realizable learnability in the well-known PAC and online classification settings. Our main result relates the evolution complexity to the DS and Littlestone dimensions, which characterize PAC and online classification respecitively.
Theorem 3.5 (Relations to the DS and Littlestone dimension).
The following statements are true.
  • (i)
    There exists $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that $\operatorname{DS}(\mathcal{F})=\infty$ but $\operatorname{C}_{T}(\mathcal{F})=\Theta(\log(T)).$
  • (ii)
    There exists $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that $\operatorname{DS}(\mathcal{F})=1$ but $\operatorname{C}_{T}(\mathcal{F})=T$.
  • (iii)
    For any $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$, we have that $\operatorname{C}_{T}(\mathcal{F})\leq\operatorname{L}(\mathcal{F})$.
  • (iv)
    There exists $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that $\operatorname{L}(\mathcal{F})=\infty$ but $\operatorname{C}_{T}(\mathcal{F})=1$.
The proof of Theorem 3.5 is in Appendix D. Parts (i) and (ii) show that the finiteness of the DSdim is neither necessary nor sufficient for learning dynamical systems in the realizable setting. On the other hand, parts (iii) and (iv) show that finite Ldim is sufficient but not necessary for learning dynamical systems in the realizable setting. Overall, learning dynamical systems is always easier than online multiclass classification, but can be both easier and harder than multiclass PAC classification. The proof of (i), (ii), and (iv) are combinatorial in nature, while the proof of (iii) involves reducing learning-to-predict in dynamical systems to online multiclass classification.
3.3 Examples
In this section, we establish the minimax rates for discrete linear systems and linear Boolean networks. The proof of Theorem 3.6 is in Appendix E.
Theorem 3.6 (Linear Systems).
  • (i)
    Let $\mathcal{X}=\mathbb{Z}^{n}$ and $r<n$. For =$\mathcal{F}=\{x\mapsto Wx\,:\,W\in\mathbb{Z}^{n\times n},\emph{\text{ rank}}(W)\leq r\}$ and $T>r$, we have $\operatorname{C}_{T}(\mathcal{F})=r+1.$
  • (ii)
    Let $\mathcal{X}=\{0,1\}^{n}$ and $T\geq n$. For =$\mathcal{F}=\{x\mapsto Wx\,(\emph{\text{mod }}2)\,:\,W\in\mathbb{Z}^{n\times n}\}$, we have $\operatorname{C}_{T}(\mathcal{F})=n.$
  • (iii)
    Let $\mathcal{X}=\{0,1\}^{n}$ and $T\geq n$. For =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{Wx>0\}\,:\,W\in\{0,1\}^{n\times n}\}$, we have $n\leq\operatorname{C}_{T}(\mathcal{F})\leq n^{2}$.
Thresholded Boolean networks have been used to model genetic regulatory dynamics (Mendoza and Alvarez-Buylla, 1998) and social networks (Kempe et al., 2003). Modulo Boolean networks have been studied by Chandrasekhar et al. (2023) in the context of stability.
4 Agnostic Learnability
4.1 Markovian Regret
Theorem 4.1 (Minimax Expected Markovian Regret).
@@ 4.1 Markovian Regret
$\max\left\{\frac{\operatorname{L}(\mathcal{F})}{18},\frac{\sqrt{T}}{16\sqrt{3}}\right\}\leq\inf_{\mathcal{A}}\,\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})\leq\operatorname{L}(\mathcal{F})+\sqrt{T\,\operatorname{L}(\mathcal{F})\log{T}}.$
One such class exhibiting the separation is the thresholds =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{x\geq a\}\,:\,a\in(0,1)\}$ used in the proof of part (iv) in Theorem 3.5. Beyond the qualitative separation of realizable and agnostic learnability under Markovian regret, we also observe a quantitative separation in terms of possible minimax rates. Recall that Theorem 3.2 shows that every rate is possible in the realizable setting. However, Theorem 4.1 shows that only two types of rates are possible in the agnostic setting: $\Theta(T)$ whenever $\operatorname{L}(\mathcal{F})=\infty$ and $\tilde{\Theta}(\sqrt{T})$ whenever $\operatorname{L}(\mathcal{F})<\infty$. More precisely, when $\sqrt{T}\geq\operatorname{L}(\mathcal{F})$, we have a lower bound of $\Omega(\sqrt{T})$ and an upper bound of $O(\sqrt{T\operatorname{L}(\mathcal{F})\log{T}})$. This raises the natural question of what the right minimax rate is. In Appendix G, we provide an evolution class and establish a lower bound of $\Omega(\sqrt{T\operatorname{L}(\mathcal{F})})$, showing that the upper bound is tight up to $\sqrt{\log{T}}$.
Our proof of the upper bound in Theorem 4.1 reduces learning dynamical systems to online multiclass classification and uses a result due to Hanneke et al. (2023). To prove the lower bound $\frac{\operatorname{L}(\mathcal{F})}{18}$, we construct a hard stream by carefully sampling a random path down a Littlestone tree of depth $\operatorname{L}(\mathcal{F})$. To prove the lower bound of $\frac{\sqrt{T}}{16\sqrt{3}}$, we construct a hard randomized stream using just two different evolution functions in $\mathcal{F}$. The full proof is in Appendix F.
4.2 Flow Regret
The necessity of the Littlestone dimension for agnostic learnability under Markovian regret is quite restrictive. For example, a simple (but unnatural) class like one-dimensional thresholds =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{x\leq a\}:a\in(0,1)\}$ has $\operatorname{L}(\mathcal{F})=\infty$ but $\operatorname{C}_{T}(\mathcal{F})=1$. The key idea in the lower bound of Theorem 4.3 is that the adversary can simulate the online multiclass classification game, where finiteness of the Littlestone dimension is necessary, by “giving up” every other round. This is possible because of the definition of Markovian regret. In particular, the evaluation of the “best-fixed evolution function in hindsight” under Markovian regret is only penalized on one-step prediction error but not on long-term consistency of the generated dynamics starting from the initial state $x_{0}$.
This motivates the following natural question. Which evolution classes $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ are agnostic learnable under Flow regret? Is the finiteness of Ldim still necessary? Theorem 4.3, whose proof can be found in Appendix H, provides partial answers to these question by bounding the minimax expected Flow regret for classes where $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|$ is uniformly bounded.
Theorem 4.3 (Minimax Expected Flow Regret).
$\frac{\operatorname{C}_{T}(\mathcal{F})}{2}\leq\inf_{\mathcal{A}}\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})\leq\operatorname{C}_{T}(\mathcal{F})+\sqrt{\operatorname{C}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{T\,K_{\mathcal{F}}}{\operatorname{C}_{T}(\mathcal{F})}\Bigl{)}}.$
Note that the upper bound becomes vacuous when $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|=\infty$, and finding a general characterization of Flow regret learnability remains an open question. Unlike agnostic learnability under Markovian regret, the finiteness of the Ldim is not necessary for agnostic learnability under Flow regret. Indeed, while the class of one-dimensional thresholds is not agnostic learnable under Markovian regret, it is agnostic learnable under Flow regret. Although Markovian learnability is not necessary for Flow regret learnability, when $K_{\mathcal{F}}<\infty$, learnability under Markovian regret is sufficient learnability under flow regret. To see this, recall that Theorem 4.1 states that learnability under Markovian regret implies $\operatorname{L}(\mathcal{F})<\infty$. Then, since part (iii) of Theorem 3.5 states $\operatorname{C}_{T}(\mathcal{F})\leq\operatorname{L}(\mathcal{F})$, we can use Theorem 4.3 to infer that $\mathcal{F}$ is also learnable under flow regret with regret $\leq\operatorname{L}(\mathcal{F})+\sqrt{\operatorname{L}(\mathcal{F})T\ln(TK_{\mathcal{F}})}$.
Theorem 4.3 shows that realizable and agnostic learnability under Flow regret are equivalent as long as $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|<\infty$. But this equivalence breaks down when the projection sizes are unbounded, as shown by Theorem 4.4.
There exists an ordered set $\mathcal{X}$ and $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that (i) $\inf_{\mathcal{A}}\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\leq 3$ but (ii) $\inf_{\mathcal{A}}\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{T}{6}$.
To prove Theorem 4.4 (see Appendix I), we construct a class $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that on large subset of states in $\mathcal{X}^{\prime}\subset\mathcal{X}$, every function $f\in\mathcal{F}$ effectively reveals its identity.
5 Discussion and Future directions
[…]
AT and US acknowledge the support of NSF via grant DMS-2413089. US also acknowledges the support of Rackham International Student Fellowship. VR acknowledges the support of the NSF GRFP. We thank Chinmaya Kaushik for pointing out a technical result, which helped us to prove part (i) Theorem 3.6.
Appendix A Proofs for Realizable learnability
Proof A.1.
@@ Appendix A Proofs for Realizable learnability
(of lower bound of Theorem 3.1). Let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ be any evolution class and $T\in\mathbb{N}$ be the time horizon. Let $\mathcal{A}$ be any randomized learner. Our goal will be to construct a hard realizable trajectory $\{x_{t}\}_{t=0}^{T}$ such that $\mathcal{A}$’s expected number of mistakes is at least $\frac{\operatorname{C}_{T}(\mathcal{F})}{2}$. Without loss of generality, suppose $d:=\operatorname{C}_{T}(\mathcal{F})>0$ as otherwise the lower bound holds trivially. Then, by definition of the Evolution complexity, there exists a trajectory tree $\mathcal{T}$ of depth $T$ shattered by $\mathcal{F}$ with branching factor at least $d$. This means that for every path $\sigma\in\{-1,1\}^{T}$ down $\mathcal{T}$, we have $\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{T}_{t}((\sigma_{<t},-1))\neq\mathcal{T}_{t}((\sigma_{<t},+1))\}\geq d$.
@@
where the first inequality follows from the fact that $\sigma_{t}\sim\{-1,1\}$.
Proof A.2.
(of upper bound of Theorem 3.1). Let $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ be any evolution function class and let $T\in\mathbb{N}$ be the time horizon. Our goal will be to construct a deterministic learner $\mathcal{A}$ such that for any realizable trajectory $\{x_{t}\}_{t=0}^{T}$, $\mathcal{A}$ makes at most $\operatorname{C}_{T}(\mathcal{F})$ mistakes. To that end, it will be useful to define an instance-dependent version of the evolution complexity.
@@
$\operatorname{C}_{T}(V,x_{0}):=\sup\{\operatorname{B}(\mathcal{T})\,\mid\,\mathcal{T}\in\mathcal{S}(V,T)\text{ and }\mathcal{T}\text{ rooted at }x_{0}\}$
denote an instance-dependent Evolution complexity of $V$, where $\mathcal{S}(V,T)$ is set of all trajectory trees of depth $T$ shattered by $V$. Note that $\operatorname{C}_{T}(V)=\sup_{x_{0}\in\mathcal{X}}\operatorname{C}_{T}(V,x_{0})$. $\mathcal{A}$ is a version-space algorithm that uses the instance-dependent Evolution complexity to make its prediction such that whenever $\mathcal{A}$ errs, the instance-dependent Evolution complexity decreases. In this way, $\operatorname{C}_{T}(V,x_{0})$ acts as a potential function. Algorithm 1 formalizes this idea.
@@
1
2Let $V_{1}=\mathcal{F}$
3for $t=1,...,T$ do
4
5      For $x\in\mathcal{X}$, define $V_{t}^{x}=\{f\in V_{t}:f(x_{t-1})=x\}.$
6      if $\{f(x_{t-1}):f\in V_{t}\}=\{x\}$ then
7             Predict $\hat{x}_{t}=x$.
8
9      else
10
11            Predict $\hat{x}_{t}\in\operatorname*{arg\,max}_{x\in\mathcal{X}}\operatorname{C}_{T-t}(V_{t}^{x},x).$
12
13      Receive $x_{t}$ and update $V_{t+1}\leftarrow V_{t}^{x_{t}}.$
14 end for
@@
We now show that Algorithm 1 makes at most $\operatorname{C}_{T}(\mathcal{F})$ mistakes on any realizable trajectory. Let $\{x_{t}\}_{t=0}^{T}$ denote the realizable trajectory to be observed. It suffices to show that
$\operatorname{C}_{T-t}(V_{t+1},x_{t})\leq\operatorname{C}_{T-t+1}(V_{t},x_{t-1})-\mathbbm{1}\{x_{t}\neq\hat{x}_{t}\}$ (3)
@@
$\sum_{t=1}^{T}\operatorname{C}_{T-t}(V_{t+1},x_{t})\leq\sum_{t=1}^{T}\operatorname{C}_{T-t+1}(V_{t},x_{t-1})-\sum_{t=1}^{T}\mathbbm{1}\{x_{t}\neq\hat{x}_{t}\}.$
@@
$\sum_{t=1}^{T}\mathbbm{1}\{x_{t}\neq\hat{x}_{t}\}$ $\leq\sum_{t=1}^{T}\operatorname{C}_{T-t+1}(V_{t},x_{t-1})-\sum_{t=1}^{T}\operatorname{C}_{T-t}(V_{t+1},x_{t})$
$=\sum_{t=1}^{T}\Bigl{(}\operatorname{C}_{T-t+1}(V_{t},x_{t-1})-\operatorname{C}_{T-t}(V_{t+1},x_{t})\Bigl{)}$
$=\operatorname{C}_{T}(V_{1},x_{0})-\operatorname{C}_{0}(V_{T+1},x_{T})\leq\operatorname{C}_{T}(\mathcal{F}).$
as needed. To prove (3), we need to consider three cases: (a) $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})>0$ and $\hat{x}_{t}\neq x_{t}$, (b) $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})>0$ and $\hat{x}_{t}=x_{t}$, and (c) $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})=0$.